OpenStudy (anonymous):

I am unsure as to writing some of the functions on here. But here goes: Simplify: 3√125(x-10)3 Both the 3's are supposed to be exponents.

OpenStudy (anonymous):

You mean this?\[\Huge \sqrt[3]{125(x-10)^3}\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Note: 125 = 5³

OpenStudy (anonymous):

Break it into \[\Huge \sqrt[3]{125} \cdot \sqrt[3]{(x-10)^3}\]

OpenStudy (anonymous):

So when simplifying it would be ^3√5^3(x-10)^3

OpenStudy (anonymous):

Ohhhh I see

OpenStudy (anonymous):

Yeah, and you can simplify it ever more.

OpenStudy (anonymous):

Is this clear?

OpenStudy (anonymous):

^3√5^3*^3√(x-10)^3 ????

OpenStudy (anonymous):

@Helenshouse It shows that you don't see clearly :)

OpenStudy (anonymous):

How are you showing ^3 ? Is there a short key to write this?

OpenStudy (anonymous):

Sigh, no but I sure am trying!

OpenStudy (anonymous):

Let me show you an example. \[\Large \sqrt[5]{a^5} = a \text{ because } a^5 = a \cdot a \cdot a \cdot a \cdot a\] Now, what is \[\huge \sqrt[3]{5^3}\] and \[\Large \sqrt[3]{(x-10)^3}\]

OpenStudy (anonymous):

The first is ^3√125 right?

OpenStudy (anonymous):

The second is ^3√(x-10)^3 I am unsure........

OpenStudy (anonymous):

No, I mean simplify it.

OpenStudy (anonymous):

@Helenshouse Let me put this way, can you convert cube root into exponent?

OpenStudy (anonymous):

Do you know any exponent properties?

OpenStudy (anonymous):

^3√125*(x-10)^3

OpenStudy (anonymous):

No, don't combine them. It makes it worse. Here, let me show you. √25 = √5² = 5, right? Do you know why?

OpenStudy (anonymous):

I am looking in my notes.

OpenStudy (anonymous):

Yes 5*5=25

OpenStudy (anonymous):

√25 = √5² = 5 Similarly, ³√ 125 = ....?

OpenStudy (anonymous):

5^3

OpenStudy (anonymous):

I am so lost here

OpenStudy (anonymous):

You missed my point, what we're looking for is ³√ 125 ( not 125)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Let try again filling the second line by following the first line: √25 = √5² = 5 Similarly, ³√ 125 = .... = ...

OpenStudy (anonymous):

√5^3 = 5*5*5 I'm not sure

OpenStudy (anonymous):

Do you understand that √5² = (5²)^( 1/2) ?

OpenStudy (anonymous):

I think it is ^3√5^3

OpenStudy (anonymous):